# On Memorization of Large Language Models in Logical Reasoning

### v1

@@ 8 Conclusion and Future Work
We propose a memorization metric $\mathsf{LiMem}$ based on the inconsistency when solving a locally perturbed logical reasoning puzzle, and quantitatively characterize the amount of memorization and reasoning. Specifically, we observe that both off-the-shelf and fine-tuned models rely on memorization to solve logical puzzles. Memorizing the training examples, however, does not seem to impede the reasoning capabilities. Through an in-depth analysis based on local perturbation, transferability, intermediate outputs probing, and fine-tuning with wrong answers, we find that LLMs learn to reason as they memorize more training examples. Furthermore, we study input and model-based signals that determine which puzzles are solved by reasoning vs by memorization. To support these studies, we create a feature-rich dynamic logical reasoning benchmark that not only enables our memorization study, but could also be useful for future studies related to LLM logical reasoning.
@@ Acknowledgments
Appendix A Details on K&K Benchmark
A.1 The Abstract Representation
@@
A.2 The Abstract Puzzle Module: Generator
@@
A.3 The Abstract Puzzle Module: Solver and Reasoner
@@
The reasoning steps are generated and stored using a similar format as the abstract representation of the puzzle as described in § A.1. The following snippet shows an example of the generated reasoning steps for the example puzzle shown above:
@@
A.4 The Abstract Puzzle Module: Perturber
@@
A.5 The Natural Language Module: NL-Generator
@@
A.6 The Natural Language Module: NL-Reasoner
@@
A.7 The Natural Language Module: NL-Perturber
@@ A.8 Dataset Generation
(a) K&K questions
(b) K&K answers
(c) 3-people K&K synthetic CoTs
(d) 5-people K&K synthetic CoTs
(e) 8-people K&K synthetic CoTs
@@
Appendix B Experimental Setups
B.1 Models
@@
B.2 Experimental Details
B.2.1 Evaluation
@@
In addition to the 0-shot direct prompting used in the main paper, we explore 0-shot Chain of Thought (CoT) prompting and 1-shot direct/CoT prompting and report the results in Appendix § C.
@@ B.2.2 Fine-tuning
For Llama3-8B fine-tuning, we used LoRA fine-tuning with the following standard hyperparameters: a batch size of 4, gradient accumulation steps of 8, and 5e-5 learning rate. We fine-tune for a maximum of 100 epochs. We primarily reported results before 50 epochs, as we found the model typically converged by then.
@@ B.2.3 Probing
B.2.4 Distinguishing Memorization from Reasoning
@@ B.2.5 Computation Resources
Appendix C Additional Experimental Results
C.1 Memorization Measurement
@@
We evaluate Llama3-8B, Phi-3-mini, Phi-3-medium, NuminaMath-7B-CoT, and Deepseek-Math-7B using 0/1-shot Direct/CoT prompting in Fig. 17. The results indicate that these open-source models exhibit poor accuracy on K&K tasks, particularly as the number of people in the K&K puzzles increases. Different prompting methods do not significantly enhance performance. Additionally, the models struggle to consistently solve the K&K prompts under local perturbations, as shown by the memorization scores under statement and lead perturbations.
@@
As shown in Fig. 18, the test accuracy (y-axis) of CoT-FTed or Direct-FTed GPT4o-mini on the unseen test set continues to increase over the epochs, despite that the memorization score $\mathsf{LiMem}(f;\mathsf{Tr})$ on training samples also increases. The memorization score $\mathsf{LiMem}(f;\mathsf{Tr})$ under role-flipping is significantly higher than other perturbation, possibly due to an internal bias that knights are truthful.
Figure 18: Test accuracy of fine-tuned GPT4o-mini (first row: Direct FT; second row: CoT FT) increase with $\mathsf{Acc}(f;\mathsf{Tr})$, despite that the memorization becomes stronger as reflected by larger $\mathsf{LiMem}(f;\mathsf{Tr})$ under leaf perturbation.
C.2 Evaluation on Reasoning Capability
C.2.1 Llama3-8B
@@ C.2.2 GPT4o-mini
C.3 Direct FT with Wrong Answers
C.3.1 Llama3-8B
@@
Figure 27: Improvement in test accuracy on $N$-person problems for Llama3-8B fine-tuned on $M$-person problems with completely wrong answers, compared to the unfine-tuned model, under various evaluation configurations.
@@
Figure 28: Improvement in test accuracy on $N$-person problems for Llama3-8B fine-tuned on $M$-person problems with $\mathbf{75\%}$ wrong answers, compared to the unfine-tuned model, under various evaluation configurations.
@@
Figure 29: Improvement in test accuracy on $N$-person problems for Llama3-8B fine-tuned on $M$-person problems with $\mathbf{50\%}$ wrong answers, compared to the unfine-tuned model, under various evaluation configurations.
C.3.2 GPT4o-mini
@@ C.4 Probing
C.5 Distinguishing Memorization from Reasoning
@@
We report test AUC for classifying puzzles based on whether they are consistently solved under leaf/statement perturbation by the Llama3-8B model Direct-FTed on the 3/5-person task. As shown in Fig. 33, the embeddings across different layers of the fine-tuned Llama3-8B provide more distinguishable signals for memorized samples than those of the base model.
(a) 3-person puzzles under leaf perturbation.
(b) 3-person puzzles under statement perturbation.
(c) 5-person puzzles under leaf perturbation.
(d) 5-person puzzles under statement perturbation.

### v2

@@ 8 Conclusion
We propose a memorization metric $\mathsf{LiMem}$ based on the inconsistency when solving a locally perturbed logical reasoning puzzle, and quantitatively characterize the amount of memorization and reasoning. Through an in-depth analysis based on local perturbation, transferability, intermediate outputs probing, and fine-tuning with wrong answers, we find that LLMs learn to reason as they memorize more training examples. Furthermore, we study input and model-based signals that may determine which puzzles are solved by reasoning vs memorization. To support these studies, we create a feature-rich dynamic logical reasoning benchmark that not only enables our memorization study, but could also be useful for future studies related to LLM logical reasoning. We defer more discussion and future work to § A.
@@ Acknowledgments
Appendix A Discussion and Future Work
Appendix B Extended Related Work
Memorization in LLMs. Prior work has explored training data memorization in LLMs, primarily in the contexts of privacy and copyright concerns (Carlini et al., 2021; Lukas et al., 2023; He et al., 2024), focusing on how LLMs may reproduce text near-verbatim to their training data (Lee et al., 2022; Carlini et al., 2023; Biderman et al., 2024). Prashanth et al. (2024) further introduces a taxonomy for memorization, categorizing it into Recitation, Reconstruction, and Recollection. They investigate the memorization behaviors of the Pythia model (Biderman et al., 2024) on the Pile dataset (Gao et al., 2020). In contrast, we examine memorization in the reasoning context, and focus on analyzing whether LLMs can accurately solve problems encountered during training but struggle to solve slightly perturbed variants. This allows us to better investigate the extent to which LLMs truly understand and generalize the underlying principles of the reasoning problems they have been trained on, as opposed to merely memorizing the text.
Recent research discusses signs of LLMs memorization in reasoning tasks by evaluating them on counterfactual reasoning tasks. These counterfactual tasks demand similar abstract reasoning skills as the original tasks but are less common in the training data. For instance, tasks such as reversing a sequence of words (McCoy et al., 2024) show better performance on high-probability sequences than on low-probability sequences; shifting each letter by $n$ places in the alphabet (Rot-$n$) (Prabhakar et al., 2024; McCoy et al., 2024) demonstrates higher performance when $n=13$ than for other values, likely because “Rot-13” is commonly used in online forums.  (Wu et al., 2024b) presents 11 counterfactual tasks (e.g., 1-indexing in Python, base-9 arithmetic) that show significant performance declines. (Jiang et al., 2024) changes some tokens in the reasoning task descriptions which leads to significant performance drops, suggesting that models might depend on recognizing superficial patterns with strong token bias. Moreover, (Razeghi et al., 2022) finds a strong correlation between the accuracy for a number on numerical reasoning tasks and its frequency in pretraining for GPT-J/GPT-Neo. In our study, we formally define a memorization score to quantify performance variance under task perturbations, covering both counterfactual alterations (e.g., switching the roles of knights and knaves) and standard perturbations on language level and problem structure level.
Detecting benchmark contamination. Recent work has shown that LLMs’ performance drastically declines when faced with altered versions of popular reasoning benchmarks, suggesting potential contamination/memorization of these benchmarks. The benchmark variants include diverse forms such as altered multiple-choice questions formats (Wang et al., 2024b; Zong et al., 2024; Gupta et al., 2024; Zhou et al., 2024; Robinson & Wingate, 2023), rephrased or translated problems (Xu et al., 2024; Yang et al., 2023; Yao et al., 2024), shuffled example orderings (Oren et al., 2024), human-curated problems of comparable difficulty (Zhang et al., 2024), functional variants generating random instantiations (Srivastava et al., 2024; Mirzadeh et al., 2024), and problems beyond specific date cutoffs (Roberts et al., 2023; Jain et al., 2024). Previous work either focus on surface level language perturbations or require extensive expert-level annotations for math level variations. In contrast, our benchmark support automatic problem-level perturbation, solution and reasoning procedure generation, and easily scale to different difficult levels and dataset sizes without extra human efforts.
Logical reasoning benchmarks. To evaluate logical reasoning capabilities in LLMs, synthetic benchmarks have been developed to enable scalable generation of samples with varying configurations and difficulty levels (Clark et al., 2020; Giadikiaroglou et al., 2024; Parmar et al., 2024). For instance, DyVal (Zhu et al., 2024) uses directed acyclic graphs to dynamically generate samples on reasoning tasks including deductive, Boolean, and abductive reasoning. (Chen et al., 2024) focus on propositional logical problems involving definite clauses, and synthetically generate variations with different premise orders, such as forward, backward, and shuffled. (Dziri et al., 2024) explore the limitations of LLMs in tasks requiring compositional reasoning, including multiplication, logic grid puzzles, and dynamic programming problems. ZebraLogic (Lin et al., 2024) is an extended benchmark that systematically tests logical reasoning capabilities. BoardgameQA (Kazemi et al., 2024) presents a question-answering dataset characterized by contradictory facts and rules in the questions. PRONTOQA (Saparov & He, 2023) is a synthetic question-answering dataset where each example is generated from a synthetic world model represented in first-order logic. This dataset enables parsing the generated chain of thought into symbolic proofs, facilitating formal analysis. TruthQuest (Mondorf & Plank, 2024) is the most similar task to our work, which provides evaluation samples based on K&K-type of puzzles involving 3-6 person. Our work provides more comprehensive dynamic set of K&K puzzles that support automatic generation of perturbations, solutions and detailed reasoning steps. Moreover, based on this benchmark, we define and measure memorization in reasoning tasks, revealing intricate interplay between memorization and reasoning in LLMs.
Improving reasoning via fine-tuning. Prior work has explored fine-tuning LLMs on synthetic reasoning data to enhance their performance on reasoning. DyVal (Zhu et al., 2024) shows that fine-tuning Llama2-13B-chat on their synthetic reasoning benchmark improves its performance on other popular reasoning benchmarks. BoardgameQA (Kazemi et al., 2024) find that fine-tuning BERT-large and T5-XXL on their training dataset with synthetic proofs outperforms few-shot CoT prompting using PaLM. (Ye et al., 2024) pretrain GPT2 from scratch on synthetic math problems, synthetic CoT steps and solutions and show that model can solve problems from the same distribution and generalize to out-of-distribution (OOD) problems. However, (Dziri et al., 2024) show that while GPT-3 fine-tuned on their compositional reasoning tasks with/without reasoning steps can solve in-distribution (ID) problems, it fails to generalize to OOD tasks with increased problem sizes. Besides using synthetic CoTs, there are work using model-generated CoTs to enhance the models’ reasoning capabilities (Chung et al., 2024). STaR (Zelikman et al., 2022) uses model self-generated CoTs on correctly solved samples to iteratively fine-tune itself as a self-taught reasoner. A number of work (Puerto et al., 2024; Kim et al., 2023; Ho et al., 2023; Hsieh et al., 2023) leverage CoTs generated from teacher models to train smaller student models. Additionally, some recent efforts have focused on leveraging intermediate reasoning steps in CoT more implicitly. For instance, Deng et al. (2023) distill intermediate reasoning tokens into the network layers by representing reasoning steps as vectors and using them as targets; Deng et al. (2024) distill CoT by gradually removing the intermediate steps and fine-tuning the model to internalize these steps, predicting the answers based on partial CoT. Both studies show that full CoT fine-tuning may not be necessary for the model to achieve strong reasoning performance.
Orthogonal to our work, inference-time techniques have been explored to enhance reasoning performance such as self-consistency (Wang et al., 2023b), self-verification (Weng et al., 2023), and integration with external symbolic solvers (Pan et al., 2023).
Grokking. Our findings are related to Grokking, first identified by  (Power et al., 2022) on a small algorithmic dataset, where validation accuracy suddenly improves from random chance to near-perfect generalization long after severe overfitting. Follow-up studies expanded the range of tasks where grokking occurs and proposed various explanations  (Liu et al., 2022a; Murty et al., 2023; Liu et al., 2022b). Recently,  (Wang et al., 2024a) observed grokking in the domain of complex knowledge-based tasks, showing that implicit reasoning over parametric knowledge emerges only after extensive overfitting. In this work, we observe a related phenomenon but through the lens of memorization and logical reasoning. Through novel (math & language-level) perturbation tests and transferability analyses, we verify that LLM’s reasoning skills emerge alongside memorization. Furthermore, our investigation focuses on logical reasoning, offering new insights into how LLMs acquire logical reasoning skills.
Adversarial robustness under perturbations. Language-level perturbations have been widely used to assess the adversarial robustness of language models, often involving manually annotated attacks, as seen in advGLUE (Wang et al., 2021), ANLI (Nie et al., 2020), and TextFooler (Jin et al., 2020). However, these approaches fundamentally differ from our proposed mathematical-level perturbations in purpose, methodology, and scope. Specifically, prior studies primarily focus on natural language understanding tasks, such as sentiment analysis and textual entailment, aiming to generate adversarial perturbations that cause misclassification without altering the ground truth. In contrast, our proposed perturbation method operates at a mathematical level in logical reasoning tasks and modifies not only the problem, but also the ground-truth answer. These mathematical perturbations ensure that the perturbed puzzle has a distinctly different solution compared to the original puzzle, while remaining superficially similar and maintaining a comparable difficulty level. This is guaranteed by the Perturber, Reasoner, and Solver components in our data generation framework. This approach provides a direct evaluation of the models’ understanding of the underlying mathematical principles. By addressing logical reasoning robustness through mathematical-level perturbations, our work contributes a novel perspective distinct from prior studies.
Appendix C Details on K&K Benchmark
C.1 The Abstract Representation
@@
C.2 The Abstract Puzzle Module: Generator
@@
C.3 The Abstract Puzzle Module: Solver and Reasoner
@@
The reasoning steps are generated and stored using a similar format as the abstract representation of the puzzle as described in § C.1. The following snippet shows an example of the generated reasoning steps for the example puzzle shown above:
@@
C.4 The Abstract Puzzle Module: Perturber
@@
C.5 The Natural Language Module: NL-Generator
@@
C.6 The Natural Language Module: NL-Reasoner
@@
C.7 The Natural Language Module: NL-Perturber
@@ C.8 Dataset Generation
Figure 14: K&K questions
Figure 15: K&K answers
Figure 16: 3-people K&K synthetic CoTs
Figure 17: 5-people K&K synthetic CoTs
Figure 18: 8-people K&K synthetic CoTs
@@
Appendix D Experimental Setups
D.1 Models
@@
D.2 Experimental Details
D.2.1 Evaluation
@@
In addition to the 0-shot direct prompting used in the main paper, we explore 0-shot Chain of Thought (CoT) prompting and 1-shot direct/CoT prompting and report the results in Appendix § E.
@@ D.2.2 Fine-tuning
For Llama3-8B fine-tuning, we used LoRA fine-tuning with the following standard hyperparameters: a batch size of 4, gradient accumulation steps of 8, and 5e-5 learning rate. The LoRA configuration was set as follows: rank $r=32$, scaling factor $\alpha=32$, and dropout rate $0.05$. No quantization techniques were used. We fine-tune for a maximum of 100 epochs. We primarily reported results before 50 epochs, as we found the model typically converged by then.
@@ D.2.3 Probing
D.2.4 Distinguishing Memorization from Reasoning
@@ D.2.5 Computation Resources
Appendix E Additional Experimental Results
E.1 Benchmark Performance of Off-the-shelf Models
@@
We evaluate Llama3-8B, Phi-3-mini, Phi-3-medium, NuminaMath-7B-CoT, and Deepseek-Math-7B using 0/1-shot Direct/CoT prompting in Fig. 21. The results indicate that these open-source models exhibit poor accuracy on K&K tasks, particularly as the number of people in the K&K puzzles increases. Different prompting methods do not significantly enhance performance.
Moreover, we evaluate GPT4o-mini under the self-consistency (Wang et al., 2023b) where we query each puzzle 40 times under temperature 0.7. Tab. 3 shows that self-consistency provides limited improvement on the 3-ppl task and fails to enhance performance on the more challenging 8-ppl task, suggesting that the model fundamentally struggles with solving such complex problems.
@@
Table 3: Self-consistency (Wang et al., 2023b) can enhance the accuracy of GPT4o-mini on the easy 2-ppl K&K task, but has limited improvement on 3-ppl task and fails to help on the challenging 8-ppl task, which suggests that the model cannot fundamentally solve such complex problems.
Method Test Accuracy
2-ppl 3-ppl 8-ppl
Direct Prompting 0.63 0.42 0.01
Direct Prompting + Self-consistency 0.74 0.43 0.02
E.2 Memorization Measurement
@@
As shown in Fig. 22, the inconsistency ratio on correctly solved training puzzles ($y$-axis) of CoT-FTed or Direct-FTed GPT4o-mini tends ot decreases over the training epochs ($x$-axis), despite that the memorization score $\mathsf{LiMem}(f;\mathsf{Tr})$ on training samples also increases (i.e., a larger proportion of memorized samples in the training set). The memorization score $\mathsf{LiMem}(f;\mathsf{Tr})$ under role-flipping is significantly higher than other perturbation, possibly due to an internal bias that knights are truthful.
Figure 22: Inconsistency ratio ($y$-axis) of fine-tuned GPT4o-mini (first row: Direct FT; second row: CoT FT) decreases over training epochs ($x$-axis), despite that the memorization becomes stronger as reflected by larger $\mathsf{LiMem}(f;\mathsf{Tr})$ (deeper color).
Figure 24: Consistency Ratio (CR $\uparrow$) under local perturbations. Fine-tuned LLMs generally demonstrate a higher consistency ratio on solved problems in the test set compared to the train set, particularly for challenging tasks such as 5/8-person puzzles. On the 3-person puzzle task, the consistency ratio between the train and test sets remains comparable. The consistency ratio generally is higher in easy tasks than in hard tasks.
Figure 25: Memorization scores of Directly Fine-Tuned Llama3-8B under various math-level (statement, leaf) and language-level (name, reorder) perturbations. Combining math-level and language-level perturbations progressively can result in higher memorization scores (e.g., leaf + reorder), especially compared to applying language-level perturbations alone.
E.3 Evaluation on Reasoning Capability
E.3.1 Llama3-8B
@@ E.3.2 GPT4o-mini
E.4 Direct FT with Wrong Answers
E.4.1 Llama3-8B
@@
Figure 34: Improvement in test accuracy on $N$-person problems for Llama3-8B Direct FTed on $M$-person problems with completely wrong answers, compared to the unfine-tuned model, under various evaluation configurations.
@@
Figure 35: Improvement in test accuracy on $N$-person problems for Llama3-8B Direct FTed on $M$-person problems with $\mathbf{75\%}$ wrong answers, compared to the unfine-tuned model, under various evaluation configurations.
@@
Figure 36: Improvement in test accuracy on $N$-person problems for Llama3-8B Direct FTed on $M$-person problems with $\mathbf{50\%}$ wrong answers, compared to the unfine-tuned model, under various evaluation configurations.
E.4.2 GPT4o-mini
@@ E.5 Probing
E.6 Distinguishing Memorization from Reasoning
@@
We report test AUC for classifying puzzles based on whether they are consistently solved under leaf/statement perturbation by the Llama3-8B model Direct-FTed on the 3/5-person task. As shown in Fig. 40 and Fig. 41, the embeddings across different layers of the fine-tuned Llama3-8B provide more distinguishable signals for memorized samples than those of the base model.
(a) 3-person puzzles under statement perturbation.
(b) 5-person puzzles under leaf perturbation.
(c) 5-person puzzles under statement perturbation.